# Adaptive Sampling for Hydrodynamic Stability

### v1

@@ 4.2 Rayleigh–Bénard convection problem
For this problem, the initial training set is generated by sampling the $(\mathrm{Ra},\delta)$ parameter domain on a uniform grid. Sixteen Rayleigh numbers are selected uniformly over the interval $\mathrm{Ra}\in[1300,1600]$, and eight perturbation amplitudes are constructed by choosing exponents uniformly in [16,2][-16,-2] and taking the corresponding powers of ten, giving $\delta\in[10^{-16},10^{-2}]$. This results in a total of $128$ parameter pairs. For each pair, the governing equations are solved using the IFISS software, and the resulting solution is labeled as bifurcated or nonbifurcated using the stability criterion described in [silvester2024machine]. The classifier retains the same architecture as in the channel-flow case and is trained for $n_{e}^{(c)}=3000$ epochs using stochastic gradient descent with batch size $1$ and learning rate $0.15$. To train the KRnet model, a set of $n_{\text{cand}}=6000$ uniformly sampled parameter points in the normalized domain is evaluated by the classifier to compute entropy-based weights. KRnet again uses $6$ affine coupling layers, and for this problem it is trained for $n_{e}^{(g)}=2000$ epochs using the WNLL loss fuction and the ADAM optimizer with batch size $500$ and learning rate $10^{-3}$. At each adaptive refinement step, KRnet contributes $n_{\text{new}}=60$ new parameter points to the training set, and two such refinements are performed for this problem.
@@ 4.3 Differentially heated cavity
For this example, the initial training set is formed by sampling eight fixed Prandtl numbers, representing a range of fluids from air ($\mathrm{Pr}=0.71$) and water ($\mathrm{Pr}=7.1$) to more viscous liquids ($\mathrm{Pr}=140,195,250,500$) and finally glycerol ($\mathrm{Pr}\approx 1000$). For each $\mathrm{Pr}$, eleven Rayleigh numbers are chosen uniformly in the interval $\mathrm{Ra}\in[2.7\times 10^{9},\,3.2\times 10^{9}]$, providing a total of $88$ initial parameter samples. Each solution is computed with IFISS and labeled according to whether the system settles to a steady state or exhibits time-periodic oscillations as in [silvester2024machine]. The classifier and KRnet architectures, as well as all training hyperparameters are taken to be the same as in problem 4.2. At each refinement step, KRnet contributes $n_{\text{new}}=50$ new parameter points, which are then added to the training set for the subsequent classifier update.
@@
Across all three problems, the adaptive sampling strategy was able to learn the bifurcation boundaries using only a small set of initially uniform samples and without any prior assumptions about the structure of the parameter space. This contrasts with the earlier study [silvester2024machine], which depended on problem-specific heuristics to construct the training set. By reducing the number of expensive IFISS simulations needed to resolve the stability transition, the proposed approach offers a flexible and efficient alternative for exploring complex fluid-dynamical bifurcations.

### v2

@@ 4.2 Rayleigh–Bénard convection problem
For this problem, the initial training set is generated by sampling the $(\mathrm{Ra},\delta)$ parameter domain on a uniform grid. Sixteen Rayleigh numbers are selected uniformly over the interval $\mathrm{Ra}\in[1300,1600]$, and eight perturbation amplitudes are constructed by choosing exponents uniformly in [16,2][-16,-2] and taking the corresponding powers of ten, giving $\delta\in[10^{-16},10^{-2}]$. This results in a total of $128$ parameter pairs. For each pair, the governing equations are solved using the IFISS software, and the resulting solution is labeled as bifurcated or nonbifurcated using the stability criterion described in [18]. The classifier retains the same architecture as in the channel-flow case and is trained for $n_{e}^{(c)}=3000$ epochs using stochastic gradient descent with batch size $1$ and learning rate $0.15$. To train the KRnet model, a set of $n_{\text{cand}}=6000$ uniformly sampled parameter points in the normalized domain is evaluated by the classifier to compute entropy-based weights. KRnet again uses $6$ affine coupling layers, and for this problem it is trained for $n_{e}^{(g)}=2000$ epochs using the WNLL loss fuction and the ADAM optimizer with batch size $500$ and learning rate $10^{-3}$. At each adaptive refinement step, KRnet contributes $n_{\text{new}}=60$ new parameter points to the training set, and two such refinements are performed for this problem.
@@ 4.3 Differentially heated cavity
For this example, the initial training set is formed by sampling eight fixed Prandtl numbers, representing a range of fluids from air ($\mathrm{Pr}=0.71$) and water ($\mathrm{Pr}=7.1$) to more viscous liquids ($\mathrm{Pr}=140,195,250,500$) and finally glycerol ($\mathrm{Pr}\approx 1000$). For each $\mathrm{Pr}$, eleven Rayleigh numbers are chosen uniformly in the interval $\mathrm{Ra}\in[2.7\times 10^{9},\,3.2\times 10^{9}]$, providing a total of $88$ initial parameter samples. Each solution is computed with IFISS and labeled according to whether the system settles to a steady state or exhibits time-periodic oscillations as in [18]. The classifier and KRnet architectures, as well as all training hyperparameters are taken to be the same as in problem 4.2. At each refinement step, KRnet contributes $n_{\text{new}}=50$ new parameter points, which are then added to the training set for the subsequent classifier update.
@@
Across all three problems, the adaptive sampling strategy was able to learn the bifurcation boundaries using only a small set of initially uniform samples and without any prior assumptions about the structure of the parameter space. This contrasts with the earlier study [18], which depended on problem-specific heuristics to construct the training set. By reducing the number of expensive IFISS simulations needed to resolve the stability transition, the proposed approach offers a flexible and efficient alternative for exploring complex fluid-dynamical bifurcations.
@@ 5 Conclusions
Reproducibility Statement
All computational results reported in this paper are reproducible. The hydrodynamic stability datasets were generated using the IFISS software package (version 3.7), which is available from https://www.manchester.ac.uk/ifiss and https://github.com/mcbssds/IFISS_download. Users must obtain IFISS separately. The adaptive sampling framework incorporates the flow-based deep generative model KRnet, obtained from the authors’ public repository (https://github.com/MJfadeaway/DAS); this code is an external dependency and is not redistributed here. Our MATLAB wrapper scripts for running IFISS in batch/parallel mode, together with the Python codes used for adaptive sampling, KR-NET integration, and surrogate model training, are available at https://github.com/SinghA01-ml/adaptive-hydrostability. Detailed installation and usage instructions are provided in the repository README.